Let x determine a random variable, and use your knowledge of probability to prepare a probability distribution: 3 cards are drawn from a deck. The number of aces is counted.
The answer to this is:
number of aces: 0,1,2,3,4
Probability: 0.7826, 0.2042, 0.0130, 0.0002
I don't understand how they arrived at this answer. Please explain.

Answer by stanbon(75887)

Let x determine a random variable, and use your knowledge of probability to prepare a probability distribution: 3 cards are drawn from a deck. The number of aces is counted.
The answer to this is:
number of aces: 0,1,2,3,4
Probability: 0.7826, 0.2042, 0.0130, 0.0002
----
There are 4 aces and 48 non-ace cards.
---
P(x = 0) = 48C3/52C3 = 0.7826
--------------------------------
P(x = 1) = [4C1*48C2]/52C3 = 4512/22100 = 0.2042
--------------------------------------
P(x = 2) = [4C2*48C1]/52C3
--------------------------------------
P(x = 3) = [4C3]/52C3
===========================================
Cheers,
Stan H.